# Overflow Prevention Enhances Long-Context Recurrent LLMs

### v1

@@ Abstract
A recent trend in LLMs is developing recurrent sub-quadratic models that improve long-context processing efficiency. We investigate leading large long-context models, focusing on how their fixed-size recurrent memory affects their performance. Our experiments reveal that, even when these models are trained for extended contexts, their use of long contexts remains underutilized. Specifically, we demonstrate that a chunk-based inference procedure, which identifies and processes only the most relevant portion of the input can mitigate recurrent memory failures and be effective for many long-context tasks: On LongBench, our method improves the overall performance of Falcon3-Mamba-Inst-7B by 14%, Falcon-Mamba-Inst-7B by 28%, RecurrentGemma-IT-9B by 50%, and RWKV6-Finch-7B by 51%. Surprisingly, this simple approach also leads to state-of-the-art results in the challenging LongBench v2 benchmark, showing competitive performance with equivalent size Transformers. Furthermore, our findings raise questions about whether recurrent models genuinely exploit long-range dependencies, as our single-chunk strategy delivers stronger performance - even in tasks that presumably require cross-context relations.
@@
(Left) We quantify this behavior by measuring zero-shot associative recall curves: The x-axis represents the number of facts (key-value pairs) in the context, while the y-axis shows the retrieval accuracy of the correct value from the context. To mitigate the memory problem, we propose OPRM, a chunk-based inference strategy that does not force the model to encode more information than it can reliably store. (Right) Leading recurrent LLMs evaluated on LongBench and LongBench_e (’0-4K’, ’4-8K’, ’8K+’). Surprisingly, our simple OPRM approach improves these models in long-context tasks.
@@ 1 Introduction
We hypothesize that the key bottleneck lies in how relevant information is captured and utilized across the entire input. While previous studies utilize toy problems such as needle in a haystack to study the limitations of recurrent models (Ben-Kish et al., 2025), here we utilize the Associative Recall (AR) task to explore their limitations. We find that even very large hidden states, such as the one of Falcon-Mamba-Inst-7B (Team, 2024), have a relatively limited memorization capacity. We quantify this phenomenon with the AR measurements, which uncover an overflow-like behavior that occurs even for short sequence inputs. This simple setting suggests that even if recurrent LLMs accept extended sequences, they do not fully leverage all the information in the context.
@@ 2 Related Work
Mamba layers are built on an evolved form of state-space layers introduced by Gu et al. (2021b; a). They exhibit promising results across several domains, including NLP (Waleffe et al., 2024; Wang et al., ), audio (Miyazaki et al., 2024), image processing (Zhu et al., 2024; Liu et al., 2024), RL (Lv et al., 2024), and more (Behrouz & Hashemi, 2024). Recent works scaled Mamba up to 7B parameters (Zuo et al., 2024; Team, 2024). The most advanced models, Falcon-Mamba and Falcon3-Mamba, were trained on 5.5T/7T tokens with sequence lengths of up to 8K/32k tokens, respectively. These models exhibit Transformer-level performance on short-sequence tasks, while enjoying the efficiency of the Mamba layer. A full definition of Mamba can be found in Appendix E.2.
@@
Hybrid models combine recurrent and attention layers to enhance both the effectiveness and efficiency of LLMs (Poli et al., 2024; Team et al., 2024b; Dong et al., 2024) demonstrating remarkable performance across multiple domains, including NLP (Ren et al., 2024; Lenz et al., 2025), vision (Hatamizadeh & Kautz, 2024), RL (Huang et al., 2024), etc. Yet, as they include attention layers, they still suffer from quadratic complexity w.r.t the input context length. Some hybrid models avoid the quadratic cost by replacing dense attention with local attention (Ren et al., 2024; Arora et al., 2024), hence enjoying sub-quadratic complexity.
RecurrentGemma. The Griffin architecture (De et al., 2024) demonstrated that a combination of gated linear recurrent units (Orvieto et al., 2023) and local attention layers can outperform SoTA Transformers at the 7B scale, while being more efficient. Scaling Griffin led to the RecurrentGemma models (Botev et al., 2024) that matched the performance of Transformer-based Gemma (Team et al., 2024a) while being trained on less tokens.
Fixed Memory Capacity in Recurrent LLMs.  Arora et al. (2024) shows that the recurrent memory capacity of the model increases as its state size grows. Guided by this principle, recent works increase the hidden state’s size in order to improve performance (Dao & Gu, 2024; Beck et al., 2024; Qin et al., 2024; Arora et al., 2024). Yang et al. (2025); Sun et al. (2025) propose recurrent layers with improved update mechanisms for better hidden state utilization. While these approaches increase memory capacity or manage it more effectively, it still remains bounded, hence may suffer from overflows. Moreover, because the capacity is fixed, re-training with a larger state may be required for certain downstream tasks.
Long-Context Capabilities. Despite their promising efficiency, SOTA recurrent and hybrid LLMs with sub-quadratic complexity have yet to match the performance of leading Transformer-based models in the regime they are designed for: real-world long-range tasks. We show that this limitation arises from their fixed memory capacity, which makes them prone to overflows, and provide a simple approach to mitigate that.
Additional related work on context-extension methods for recurrent LLMs such as DeciMamba (Ben-Kish et al., 2025) and others (Ye et al., 2025), is provided in Appendix E.
@@ 3 Problem Investigation
We use the Associative Recall task (described below) to investigate the memory overflow phenomenon in large recurrent LLMs. Memory overflows occur when a model fails to store or retrieve relevant information from the context. In the AR task, these overflows are indicated by a decrease in accuracy compared to the initial performance.
@@
Associative Recall (AR). In psychology, Associative Recall (AR) is defined as the ability to learn and remember the relationship between unrelated items. In the context of LLMs, AR plays an important role in the emergence of critical learning capabilities (Olsson et al., 2022), hence widely studied. The AR task that we use in our experiments is similar to Arora et al. (2023): The context is a concatenation of $M$ facts, which are key-value pairs $(K_{i},V_{i}),\,i\in[M]$, and are eventually followed by a query $Q$, which is equal to one of the keys in the context:
@@
The keys and values are drawn from a token vocabulary $\mathbb{V}$. In our experiments each key or value is between 1 to 3 tokens long, depending on the experiment. The objective of the model is to predict the corresponding value for the query $Q$. It is important to note that there are no contradictions between the facts, i.e. a key $K_{i}$ cannot be attributed to two different values. Lastly, to control for the effects of varying context lengths, all contexts are padded to a pre-defined length, N, using special pad tokens that are inserted between key-value pairs. Thus, all samples have the same length, regardless of the amount of facts in the context.
Zero-Shot AR with Recurrent LLMs. We evaluate a recurrent model with $M\in[15,160]$ facts, where each context is padded to a length of N=1200 tokens. Implementation details are provided in Appendix A.1. Our results are displayed in Figure 1 (left, blue curve). The graph shows the behavior of a popular recurrent LLM, Falcon-Mamba-Inst-7B, which has a relatively large hidden state size: 4096 channels, each with a state size of 16. The AR profile behaves as follows: When the number of facts is small, we observe a healthy memory behavior, achieving slightly higher than 80% accuracy. Then, as we increase the amount of facts in the context, the performance starts to drop, revealing the first occurrences of overflows. When the amount of information is increased further, the model remembers only a few single facts out of the hundreds in the context. This measurement is performed with short contexts (N=1200), emphasizing the severity of this issue in the long-context domain.
Figure 2: Memory overflows in a controlled setup. (Left) AR curves of 2-layer models with different hidden state sizes. $d$ is the channels dimension and $d_{state}$ is the state size. The x-axis represents the number of facts in the context, and the y-axis shows the retrieval accuracy of correct values from the context. (Right) Memory capacity as a function of channel and state dimensions. Each point shows the ratio between the capacity (maximal amount of retrieved facts from a single context) and the amount of facts that the model was trained to retrieve (128 facts). All 2-layer models were trained to retrieve up to 128 facts, yet most models suffer from overflows, therefore have a lower capacity than the one trained for.
AR in a Controlled Setup. To strengthen the validity of this result, we capture the overflow phenomenon in an additional independent, controlled, setup where we train 2-layer Mamba models from scratch on a synthetic AR task. We test different combinations of channels dimension $d\in\{64,128,256,512\}$ and state dimension $d_{state}\in\{1,2,4,8,16\}$, resulting in a wide variety of hidden state sizes (In Mamba, defined by $d\times d_{state}$). In addition, we use contexts of length N=384 tokens, and train each model to retrieve up to 128 facts. More implementation details are provided in Appendix A.2. As can be seen in Figure 2 (left), the 2-layer models produce an associative recall profile similar to that produced by the recurrent LLM in the zero-shot setting in Figure 1 (left, blue curve). Furthermore, when increasing the hidden state size, the overflow phenomenon becomes less severe, yet, it is not fully prevented. This highlights the problem with the ’fixed memory scaling’ approach discussed in Section 2: vanilla recurrent models will never be able to process contexts of arbitrary data content, such as large code repositories or long videos. We visualize this in Figure 2 (right), where each data point shows the memory capacity for a given hidden state size. While the capacity increases with state size, the models are still not able to retrieve all 128 facts from the context, despite being trained to do so. As for the scaling trends, we find that both channel and state dimensions are necessary for increasing the capacity - yet the amount of channels must be significantly larger than the amount of states - here ×$\times 32$.
4 Method
From Section 3, we conclude that for given channel and state dimensions, the maximum capacity of the recurrent memory is limited. Furthermore, the performance of the model degrades as the amount of information increases beyond this limit. Motivated by this observation, we propose OPRM (Overflow Prevention for Recurrent Models), a simple yet effective inference method. The core idea behind our method is to chunk the context such that the information content of each chunk does not exceed the model’s limit. Surprisingly, we find that decoding a single relevant chunk leads to significant gains over the baseline method in a variety of tasks.
Overall Mechanism. We consider a generation task where a prompt $X=[P,C,S]$, consists of a prefix $P$, a context $C$, and a suffix $S$, which contains a query $Q$. The model is then given $X$ to generate an answer $A$. Our method operates in two stages: speculative prefill and selective decoding, which are described in the next two subsections (Secs. 4.1 and 4.2). This design follows the standard prefill-decode framework in LLMs, as visualized in Figure 3. Our method is grounded in several key design principles, including the role of locality and compression in NLP, as well as a speculative processing strategy. Further details on these aspects are provided in Appendix C.
@@ 4.1 Speculative Prefill
OPRM first splits the context C into b chunks $\{C_{i}\}_{i=1}^{b}$ of the same length $L$ and constructs b separate prompts, each maintaining the original structure: $X_{i}=[P,C_{i},S]$. These prompts are processed in parallel in a speculative manner, efficiently computing the output distribution for the first answer token $Pr(\cdot\mid X_{i})$, the state after processing $X_{i}$ (denoted as $H_{i}$), and the token predicted based on $Pr(\cdot\mid X_{i})$ (denoted as $A_{i,0}$).
@@ 4.2 Selective Decoding
$j=\arg\min_{i}\{E_{i}\mid i\in[b]\},\quad E_{i}=\sum_{v\in V}Pr(v\mid X_{i})\cdot\log_{2}Pr(v\mid X_{i}),$
where $V$ represents the vocabulary, and $E_{i}$ serves as an uncertainty score for each prompt.
@@ 5 Experiments
We evaluate OPRM as an enhancer of long-context abilities of recurrent LLMs across multiple tasks, detailed in Sections 5.1 and 5.2. We then conduct ablation studies in Section 5.3 and conclude with an empirical analysis of our method’s efficiency in Section B.1 (Appendix). In all experiments we perform OPRM inference with the min entropy selection criterion. All implementation details are reported in Appendix A.

### v2

@@ Abstract
A recent trend in LLMs is developing recurrent sub-quadratic models that improve long-context processing efficiency. We investigate leading large long-context models, focusing on how their fixed-size recurrent memory affects their performance. Our experiments reveal that, even when these models are trained for extended lengths, their use of long contexts remains underutilized. Specifically, we demonstrate that a chunk-based inference procedure, which identifies and processes only the most relevant portion of the input can mitigate recurrent memory failures and be effective for many long-context tasks: On LongBench, our method improves the overall performance of Falcon3-Mamba-Inst-7B by 14%, Falcon-Mamba-Inst-7B by 28%, RecurrentGemma-IT-9B by 50%, and RWKV6-Finch-7B by 51%. Surprisingly, this simple approach also leads to state-of-the-art results in the challenging LongBench v2 benchmark, showing competitive performance with equivalent size Transformers. Furthermore, our findings raise questions about whether recurrent models genuinely exploit long-range dependencies, as our single-chunk strategy delivers stronger performance - even in tasks that presumably require cross-context relations.
@@
(Left) We quantify this behavior by measuring zero-shot associative recall curves: The x-axis represents the number of facts (key-value pairs) in the context, while the y-axis shows the retrieval accuracy of the correct value from the context. To mitigate the memory problem, we propose OPRM, a chunk-based inference strategy that does not force the model to encode more information than it can reliably store. (Right) Leading recurrent LLMs evaluated on LongBench and LongBench_e (’0-4K’, ’4-8K’, ’8K+’). Surprisingly, our simple approach significantly improves performance on long-context tasks.
@@ 1 Introduction
We hypothesize that the key bottleneck lies in how relevant information is captured and utilized across the entire input. While previous studies utilize toy problems such as needle in a haystack to study the limitations of recurrent models (Ben-Kish et al., 2025), here we utilize the Associative Recall (AR) task to explore their limitations. We find that even very large hidden states, such as the one of Falcon-Mamba-Inst-7B (Zuo et al., 2024), have a relatively limited memorization capacity. We quantify this phenomenon with the AR measurements, which uncover an overflow-like behavior that occurs even for short sequence inputs. This simple setting suggests that even if recurrent LLMs accept extended sequences, they do not fully leverage all the information in the context.
@@ 2 Related Work
Mamba layers are built on an evolved form of state-space layers introduced by Gu et al. (2021b; a). They exhibit promising results across several domains, including NLP (Waleffe et al., 2024; Wang et al., ), audio (Miyazaki et al., 2024), image processing (Zhu et al., 2024; Liu et al., 2024b), RL (Lv et al., 2024), and more (Behrouz & Hashemi, 2024). Recent works scaled Mamba up to 7B parameters (Zuo et al., 2024; Team, 2024). The most advanced model, Falcon3-Mamba, was trained on 7T tokens with sequence lengths of up to 32k tokens. These models exhibit Transformer-level performance on short-sequence tasks, while enjoying the efficiency of the Mamba layer. A full definition of Mamba can be found in Appendix E.3.
@@
Hybrid models combine recurrent and attention layers to enhance both the effectiveness and efficiency of LLMs (Poli et al., 2024; Team et al., 2024b; Dong et al., 2024) demonstrating remarkable performance across multiple domains, including NLP (Ren et al., 2024; Lenz et al., 2025), vision (Hatamizadeh & Kautz, 2024), RL (Huang et al., 2024), etc. Yet, as they include attention layers, they still suffer from quadratic complexity w.r.t the input context length. Some hybrid models avoid the quadratic cost by replacing dense attention with local attention (Ren et al., 2024; Arora et al., 2024), hence enjoying sub-quadratic complexity. The Griffin architecture (De et al., 2024) demonstrated that a combination of gated linear recurrent units (Orvieto et al., 2023) and local attention layers can outperform SoTA Transformers at the 7B scale, while being more efficient. Scaling Griffin led to the RecurrentGemma models (Botev et al., 2024) that matched the performance of Transformer-based Gemma (Team et al., 2024a) while being trained on less tokens.
Fixed Memory Capacity in Recurrent LLMs.  Arora et al. (2024) shows that the recurrent memory capacity of the model increases as its state size grows. Guided by this principle, recent works increase the hidden state’s size in order to improve performance (Dao & Gu, 2024; Beck et al., 2024; Qin et al., 2024b; Arora et al., 2024). Yang et al. (2025); Sun et al. (2025) propose recurrent layers with improved update mechanisms for better hidden state utilization. While these approaches increase memory capacity or manage it more effectively, it still remains bounded, hence may suffer from overflows. Moreover, because the capacity is fixed, re-training with a larger state may be required for certain downstream tasks.
Long-Context Capabilities. Despite their promising efficiency, SOTA recurrent and hybrid LLMs with sub-quadratic complexity have yet to match the performance of leading Transformer-based models in the regime they are designed for: real-world long-range tasks. We show that this limitation arises from their fixed memory capacity, which makes them prone to overflows, and provide a simple approach to mitigate this behavior.
Additional related work on RAG-based methods, as well as context-extension methods for recurrent LLMs such as DeciMamba (Ben-Kish et al., 2025), is provided in Appendix E.
@@ 3 Problem Investigation
We use the AR task (described below) to investigate the memory overflow phenomenon in SoTA recurrent LLMs. Memory overflows are a failure mode where the model cannot store or retrieve relevant information from the context. Specifically, when relevant information is combined with only a small amount of additional information (e.g., a few extra facts), the model can retain and retrieve it successfully. However, as more additional information is introduced, the effective state capacity is exceeded, resulting in retrieval failures. In the AR task, overflows are indicated by a drop in accuracy relative to the initial value.
@@
Associative Recall (AR). In psychology, Associative Recall (AR) is defined as the ability to learn and remember the relationship between unrelated items. In the context of LLMs, AR plays an important role in the emergence of critical learning capabilities (Olsson et al., 2022), hence widely studied (Arora et al., 2025; Lutati et al., 2023; Okpekpe & Orvieto, 2025). The AR task that we use in our experiments is similar to Arora et al. (2023): The context is a concatenation of $M$ facts, which are key-value pairs $(K_{i},V_{i}),\,i\in[M]$, and are eventually followed by a query $Q$, which is equal to one of the keys in the context.
$\text{AR Sample:}\quad{\color[rgb]{0,0,0}K_{1}\,\,V_{1}}\,\,\,{\color[rgb]{.5,.5,.5}<pad>\,\,\,<pad>}\,\,\,{\color[rgb]{0,0,0}K_{2}\,\,V_{2}}\,\,\,{\color[rgb]{.5,.5,.5}<pad>\,\,\,<pad>}\,\,\,\cdot\cdot\cdot\,\,\,{\color[rgb]{0,0,0}K_{M}\,\,V_{M}}\,\,\,{\color[rgb]{0,0,0}Q}\,\,\,{\color[rgb]{0,0,0}\textbf{?}}$
The keys and values are drawn from a token vocabulary $\mathbb{V}$. In our experiments each key or value is between 1 to 3 tokens long, depending on the experiment. The objective of the model is to predict the corresponding value for the query $Q$. It is important to note that there are no contradictions between the facts, i.e. a key $K_{i}$ cannot be attributed to two different values. Lastly, to control for the effects of varying context lengths, all contexts are padded to a pre-defined length $N$, using special pad tokens that are inserted between key-value pairs. Thus, all samples have the same length, regardless of the amount of facts in the context.
Zero-Shot AR with SoTA Recurrent LLMs. We evaluate a recurrent model with $M\in[15,160]$ facts, where each context is padded to a length of $N=1200$ tokens. Implementation details are provided in Appendix A.1. Our results are displayed in Figure 1 (left, blue curve). The graph shows the behavior of a popular recurrent LLM, Falcon-Mamba-Inst-7B, which has a relatively large hidden state size: 4096 channels, each with a state size of 16. The AR profile behaves as follows: When the number of facts is small, we observe a healthy memory behavior, achieving slightly higher than 80% accuracy. Then, as we increase the amount of facts in the context, the performance starts to drop, revealing the first occurrences of overflows. When the amount of information is increased further, the model remembers only a few single facts out of the hundreds in the context. This measurement is performed with short contexts ($N=1200$), emphasizing the severity of this issue in the long-context domain. In Section B.5 (Appendix), we measure the sensitivity of AR performance to the input sequence length $N$. We find that the model is not particularly sensitive to the context’s length, but rather to the amount of information it contains. This further shows that overflows are a fundamentally distinct limitation of recurrent LLMs - one that should be addressed independently of other known limitations, such as length generalization. Lastly, Section B.6 (Appendix) presents an analysis of how overflows vary with token position.
Figure 2: Memory overflows in a controlled setup. (Left) AR curves of 2-layer models with different hidden state sizes. $d$ is the channels dimension and $d_{state}$ is the state size. The x-axis represents the number of facts in the context, and the y-axis shows the retrieval accuracy of correct values from the context. (Right) Memory capacity as a function of channel and state dimensions. Each point shows the ratio between the capacity (maximal amount of retrieved facts from a single context) and the amount of facts that the model was trained to retrieve (here, 128 facts). All 2-layer models were trained to retrieve up to 128 facts, yet most models suffer from overflows, therefore have a lower capacity than the one trained for.
AR in a Controlled Setup. To strengthen the validity of this result, we capture the overflow phenomenon in an additional independent, controlled, setup where we train 2-layer Mamba models from scratch on a synthetic AR task. We test different combinations of channels dimension $d\in\{64,128,256,512\}$ and state dimension $d_{state}\in\{1,2,4,8,16\}$, resulting in a wide variety of hidden state sizes (In Mamba, defined by $d\times d_{state}$). In addition, we use contexts of length N=384 tokens, and train each model to retrieve up to 128 facts. More implementation details are provided in Appendix A.2. As can be seen in Figure 2 (left), the 2-layer models produce an associative recall profile similar to that produced by the SoTA recurrent LLM in the zero-shot setting in Figure 1 (left, blue curve). Furthermore, when increasing the hidden state size, the overflow phenomenon becomes less severe, yet, it is not fully prevented. This highlights the problem with the ’fixed memory scaling’ approach discussed in Section 2: vanilla recurrent models will never be able to process contexts of arbitrary data content, such as large code repositories or long videos. We visualize this in Figure 2 (right), where each data point shows the memory capacity for a given hidden state size. While the capacity increases with state size, the models are still not able to retrieve all 128 facts from the context, despite being trained to do so. As for the scaling trends, we find that both channel and state dimensions are necessary for increasing the capacity - yet the amount of channels must be significantly larger than the amount of states - here $\times 32$.
4 Method
From Section 3, we conclude that for given channel and state dimensions, the maximum capacity of the recurrent memory is limited. Furthermore, the performance of the model degrades as the amount of information increases beyond this limit. Motivated by this observation, we propose OPRM (Overflow Prevention for Recurrent Models), a simple yet effective inference method. The core idea is to chunk the context such that the information content of each chunk does not exceed the model’s limit. Surprisingly, we find that decoding a single relevant chunk yields significant gains over the baseline across a variety of tasks.
Overall Mechanism. We consider a generation task where a prompt $X=[P,C,S]$, consists of a prefix $P$, a context $C$, and a suffix $S$, which contains a query $Q$. The model is then given $X$ to generate an answer $A$. Our method operates in two stages: speculative prefill and selective decoding, which are described in the next two subsections (Secs. 4.1 and 4.2). This design follows the standard prefill-decode framework in LLMs, as visualized in Figure 3, and formally described in Algorithm 1. Our method is grounded in several key design principles, including the role of locality and compression in NLP, as well as a speculative processing strategy. Further details on these aspects are provided in Appendix C.
Algorithm 1 Overflow Prevention for Recurrent Models (OPRM)
@@ 4 Method
[…]
1. Preprocessing - prepare OPRM chunks:
 1.1 $\text{Right pad }C\text{ s.t. }|C|\bmod L=0$; $b\leftarrow|C|/L$
 1.2 $\forall i\in[b]\text{: }\ C_{i}\leftarrow C[(i-1)\!\cdot\!L:i\!\cdot\!L]$
 1.3 $\forall i\in[b]\text{: }\ X_{i}\leftarrow[P,C_{i},S]$
2. Speculative Pre-Fill:
 2.1 $\forall i\in[b]\text{ (in parallel): }\ H_{i},\,Pr(\cdot\mid X_{i})\leftarrow\mathcal{M}(0,X_{i})$ $\triangleright$ Compute state and logits for all chunks
3. Selective Decoding:
 3.1 $\Gamma_{IDK}\leftarrow\{\arg\max\{\,Pr(v\mid X_{i})\mid v\in\mathcal{V}\,\}=\text{error\_token\_id}\mid\forall i\in[b]\}$ $\triangleright$ Find IDK Chunks
 3.2 $j\leftarrow\arg\min\{\,E_{i}\mid i\in[b]/\Gamma_{IDK}\,\}$ $\triangleright$ Select decoding state (here - entropy criteria)
 3.3 $A_{j,0}\leftarrow\tau(Pr(\cdot\mid X_{j}))$ $\triangleright$ Obtain first answer token
 3.4 $A\leftarrow\tau(\mathcal{M}(H_{j},A_{j,0}))$ $\triangleright$ Decode answer AA autoregressively starting from state $H_{j}$
return AA
@@ 4.1 Speculative Prefill
OPRM first splits the context C into b chunks $\{C_{i}\}_{i=1}^{b}$ of the same length $L$ and constructs b separate prompts, each maintaining the original structure: $X_{i}=[P,C_{i},S]$. These prompts are processed in parallel in a speculative manner, efficiently computing the output distribution for the first answer token $Pr(\cdot\mid X_{i})$, the state after processing $X_{i}$ (denoted as $H_{i}$), and the token predicted based on $Pr(\cdot\mid X_{i})$ and decoding algorithm $\tau$ (denoted as $A_{i,0}$).
@@ 4.2 Selective Decoding
$j=\arg\min_{i}\{E_{i}\mid i\in[b]\},\quad E_{i}=\sum_{v\in\mathcal{V}}Pr(v\mid X_{i})\cdot\log_{2}Pr(v\mid X_{i}),$
where $\mathcal{V}$ represents the vocabulary, and $E_{i}$ serves as an uncertainty score for each prompt.
@@ 5 Experiments
We evaluate OPRM as an enhancer of long-context abilities of recurrent LLMs across multiple tasks, detailed in Sections 5.1 and 5.2. We then conduct ablation studies in Section 5.3, followed by an empirical analysis of our method’s efficiency in Section 5.4. Additional experiments are provided in Section B (Appendix): comparisons to agentic and RAG-based long-context methods, more long context benchmarks (Zhang et al., 2024), and ablations exploring multi-chunk aggregation. In all experiments we perform OPRM inference with the min entropy selection criterion. All implementation details are reported in Appendix A.